A light source moves toward a receiver with speed v. Part A If the source emits light with a frequency f0, what is the frequency of the received signals? Hint: Remember that λf=c for electromagnetic waves. Express your answer in terms of the variables f0, v, and c. f =

Part B Evaluate numerically for a source of 250 nm ultraviolet light moving toward an observer at 0.4 c.
Answer #1

the frequency of the received signals:

f = sqrt[(c + v) / (c- v)]*f0

here, f0 = c/λ

---------------------------------------------------

b.

f0 = c/λ = 3*108 / 250*10-9 = 1.2e+15 Hz

f = sqrt[(c + v) / (c- v)]*f0

= sqrt[(c+0.4c) / (c- 0.4c)]*f0

= sqrt[(1.4) / (0.6)]*1.2e+15

= 1.833e+15 Hz